C=500+0.4yd, mps=0.6,I=500,G=700,*200 and m=50+0.2Y the tax rate is 20% of Y the population is 90 people. Calculate national income equilibrium


1
Expert's answer
2021-09-22T07:27:14-0400

Solution:

Equilibrium national income:

At equilibrium: Y = AE

Y = C + G + I

C = 500 + 0.4Yd = 500 + 0.4(Y – T)

T = 20% of income (Y) = 0.2Y

C = 500 + 0.4(Y – 0.2Y)

G = 700

I = 500

Y = 500 + 0.4(Y – 0.2Y) + 700 + 500

Y = 500 + 0.32Y + 700 + 500

Y – 0.32Y = 1,700

0.68Y = 1.,600

Y = 2,500

 

The equilibrium national income = 2,500